OpenStudy (anonymous):

Evaluate -2a^2 if a = -5/2

mathslover (mathslover):

ok so can u tell me : what is \[\huge{(\frac{-5}{2})^2}\]

OpenStudy (anonymous):

6.25

OpenStudy (anonymous):

-6.25

mathslover (mathslover):

ok just tell me the value in fractions

OpenStudy (anonymous):

6 1/4

mathslover (mathslover):

ok so according to you : \[\huge{\frac{-5^2}{2^2}=\frac{61}{4}}\]

mathslover (mathslover):

so what is (-5)^2 ?

OpenStudy (anonymous):

-25/2

mathslover (mathslover):

sorry to say champs but u r wrong ..

OpenStudy (anonymous):

:O why?

OpenStudy (anonymous):

i'm lost..

OpenStudy (hba):